OpenStudy (anonymous):

Find all intercepts and asymptotes y= x^2-4/(x^2-x-12)

OpenStudy (anonymous):

For the vertical asymptotes, you need to find the values of x for which the denominator is zero\[x^2-x-12=0\]Solve this quadratic equation to get two values of x. Are you familiar with the kind?

OpenStudy (anonymous):

so X cannot be -4 and positive 3?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Wait isn't it opposite now that I think about it.

OpenStudy (anonymous):

(x+3)(x-4) = (x^2-x-12)?

OpenStudy (anonymous):

Right, (x+3)(x-4) is in the denominator and if x = -3 or x = 4, then y = infinity. So x = -3 and x = 4 are the asymptotes.